# Balanced Spanning Tree Distributions Have Separation Fairness

### v1

@@ Abstract
Sampling-based methods such as ReCom are widely used to audit redistricting plans for fairness, with the balanced spanning tree distribution playing a central role since it favors compact, contiguous, and population-balanced districts. However, whether such samples are truly representative or exhibit hidden biases remains an open question. In this work, we introduce the notion of separation fairness, which asks whether adjacent geographic units are separated with at most a constant probability (bounded away from one) in sampled redistricting plans. Focusing on grid graphs and two-district partitions, we prove that the smooth variant of the balanced spanning tree distribution satisfies separation fairness. Our results also provide theoretical support for popular MCMC methods like ReCom, suggesting that they maintain fairness at a granular level in the sampling process. Along the way, we develop tools for analyzing loop-erased random walks and partitions that may be of independent interest.
@@ 1.2 Separation Fairness and Our Result
Finally, we show that our results also imply constant separation fairness for the standard spanning tree distribution restricted to exactly balanced partitions, or to partitions where the imbalance is an additive 2, under some conjectures on the properties of the spanning tree distribution. We discuss this further in Section 6.
@@
Similarly, though the spanning tree distribution has a close connection to random walks on the underlying graph, a significant amount of the randomness is “lost” when constraints such as balance are added. For example, consider the probability that two vertices are separated in the spanning tree distribution $\mathcal{D}_{\mathtt{sp}}$. For $k=2$, this probability can be shown to be directly proportional to the effective resistance between the two vertices wu15-sepprob ; wu23-sepprob . On the other hand, no general expression is known for this probability even under mild modifications of the distribution towards balance.
@@
Definition 9 and Theorem 4 form the heart of our technical contributions. The former and the associated Theorem 5 shows that our modified process indeed allows us to analyze partitions purely from the perspective of random cycles. Our main contribution here is a constructive mapping between random walks: given a walk whose erasure separates two vertices, we devise an intricate algorithm that builds a corresponding walk that does not, carefully re-splicing all the original loops to ensure the mapping is one-to-one, and the number of different directed edges is small. This formalizes the intuition that a local change to a partition’s boundary should correspond to a local change in the underlying random walk. The latter, Theorem 4, proves that such a local modification is always possible. To do this, we develop a suite of graph-theoretic tools in Section 5.2 specifically for analyzing 2-partitions on grid graphs. Our key insight is the following: After switching vertices from one part to the other, either the original partitions were both “thick” and this did not disconnect any part, or they were “thin” and there must exist some structures that we can identify and perform local surgery on (Lemmas 22 and 23). Through an extensive case analysis facilitated by these tools, we prove that any separating partition can be modified in a small neighborhood to become non-separating, with only a constant change to the imbalance.
@@ 3.1 Modified Random Walk Algorithm
where the sum is over all connected partitions $X,Y$. Similarly, for any pair of vertices $u,v\in V(G)$, let $\texttt{sp}(G,2)_{u\not\sim v}$ denote the same sum except the sum is over all connected partitions $X,Y$ such that $u\in X$, $v\in Y$. We note that these normalization factors do not need to be computed, since we do not intend to actually run our algorithm, only analyze it. We are now ready to state Algorithm 2.
@@ 3.3 Unseparating Mappings imply Theorem 1
We assume Theorems 4 and 5 are true. Claim 2 implies that that we can sample from $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ as follows: First, run Algorithm 2. Accept the resulting partition $P$ with probability $e^{-\lambda\cdot\texttt{imb}(P)}$ and restart otherwise. Fix some pair of adjacent vertices $u,v$ and let $f$ be the promised unseparating mapping for them. Let $H^{*}$ be the subgrid in Property 1.
@@
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}\wedge\mathbf{E}_{2\beta}]\geq\frac{1}{3\beta^{2}e^{\delta\lambda}4^{3\beta^{2}}}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}\wedge\mathbf{E}_{2\beta}].$ (4)
@@
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}\wedge\mathbf{E}_{2\beta}]$ $\geq\frac{1}{3\beta^{2}e^{\delta\lambda}4^{3\beta^{2}}}(\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}]-\frac{1}{2})$
$\implies\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}]$ $\geq\frac{1}{3\beta^{2}e^{\delta\lambda}4^{3\beta^{2}}}(\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}]-\frac{1}{2})$
@@ 4.2 Proof of Theorem 5
Now, recall that Algorithm 3 added these walks to $W^{\prime}$ in such a way that $N_{1}$ appears first in $\widehat{W}$, $N_{2}$ appears just after that, and so on, so that $N_{p}$ appears last. Let $L_{1},L_{2},\ldots,L_{m^{\prime}}$ be the loops added because of $N_{1}$. That is, $N_{1}=L_{1}L_{2}\ldots L_{m^{\prime}}$. We want to show that in the first iteration of step 12 in Algorithm 4, we correctly identified $N_{1}$. For any $1\leq p\leq m^{\prime}$, let $w_{p}$ be the smallest element in $L_{p}$ according to $\preceq_{D}$, so that $\{w_{1},w_{2},\ldots,w_{m^{\prime}}\}$ is the loop sequence of these loops. Since each $L_{p}$ (or possibly a cyclic permutation of it) was a part of $M_{1}$, which was a maximal loop at $v_{i}$ in $W$, we can apply Observation 9 to get that either $w_{p}=\varphi$ or $v_{1}\preceq_{D}w_{p}$. Further, since $N_{1}$ must be the cyclic permutation of $M_{1}$ from the first appearance of $v_{1}$, $v_{1}$ must appear on $L_{m^{\prime}}$. If $v_{1}$ did not appear on $L_{m^{\prime}}$ but only appeared on, say, $L_{m^{\prime}-1}$ instead, this would mean that $N_{1}$ is the cyclic permutation of $M_{1}$ from the second appearance of $u_{i}$. Thus, $w_{m^{\prime}}=v_{1}$. We now argue that $m^{\prime}$ is the first element of the breakdown of the loop-sequence of $\mathcal{L}$. This will complete our argument that $N_{1}$ was correctly identified in step 12. To see this, observe that by construction of $\mathcal{M}_{u_{i}}$, $v_{1}\prec_{D}v_{2}\prec_{D}\ldots\prec_{D}v_{m}$. Further, the walks $M_{2},M_{3},\ldots,$ were comprised of maximal loops from $v_{2},v_{3},\ldots$ respectively. From Observation 9, for any $p>m^{\prime}$, it must be that $w_{m^{\prime}}=v_{1}\prec_{D}w_{p}$, or $w_{p}=\varphi$. Combined with our earlier observation that $0<p<m^{\prime}\implies v_{1}\preceq_{D}w_{p}$, or $w_{p}=\varphi$, this means that $v_{1}$ is in the breakdown of the loop-sequence of $\mathcal{L}$. It also must be the first element of the breakdown, since if there exists some $w_{p}$ with $p<m^{\prime}$ and $w_{p}\prec_{D}v_{1}$, this creates a contradiction with our observations. Repeating the arguments for $N_{2},N_{3},\ldots,$ in the same way completes the proof of Subclaim 0.1, when $\mathcal{M}_{u_{i^{\prime}}}=\phi$ for all $i^{\prime}<i$.
@@
This concludes the proof of Lemma 14. Along with Claims 11 and 13, this also completes the proof of Theorem 5.

### v2

@@ Abstract
Sampling-based methods such as ReCom are widely used to audit redistricting plans for fairness, with the balanced spanning tree distribution playing a central role since it favors compact, contiguous, and population-balanced districts. However, whether such samples are truly representative or exhibit hidden biases remains an open question. In this work, we introduce the notion of separation fairness, which asks whether adjacent geographic units are separated with at most a constant probability (bounded away from one) in sampled redistricting plans. Focusing on grid graphs and two-district partitions, we prove that a smooth variant of the balanced spanning tree distribution satisfies separation fairness. Our results also provide theoretical support for popular MCMC methods like ReCom, suggesting that they maintain fairness at a granular level in the sampling process. Along the way, we prove a novel local-interchangeability lemma for 2-partitions on grids, showing that any separation of adjacent vertices can be undone via a constant-sized modification. This lemma, along with our other tools for analyzing the structure of partitions and loop-erased random walks, may be of independent interest.
@@ 1.2 Separation Fairness and Our Result
The proof of Theorem 1 hinges on the following local-interchangeability result which is of interest on its own. Such a result is non-trivial even for 2-partitions on a grid. Our proof requires a careful case-analysis, aided by tools we develop in Section 5.2 that are essential in pruning the number of cases to a tractable level.
Theorem 2 (Informal version of Theorem 21).
Given any 2-partition $P$ of a $n\times m$ grid (with $n,m=\Omega(1)$), and two adjacent vertices $u,v$ that are in different parts in $P$, there is a partition $P^{\prime}$ such that $u,v$ are in the same part in $P^{\prime}$, and $P,P^{\prime}$ differ only in a constant-sized neighborhood around $u,v$.
Our results also have implications for groups of such separation events. We show that the probability that any $q$ sufficiently separated edges are simultaneously separated decays rapidly as $q$ increases.
Corollary 3 (Informal version of Theorem 9).
For any set of $q$ sufficiently separated edges of a $n\times m$ grid (with $n,m=\Omega(1)$), the probability that a 2-partition sampled from $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ separates all the $q$ edges is at most $e^{-\Omega(q)}$.
This shows a strong independence property for separation events in spanning tree distributions.
@@
Finally, we show that our results also imply constant separation fairness for the standard spanning tree distribution restricted to exactly balanced partitions, or to partitions where the imbalance is an additive 3, under some conjectures on the properties of the spanning tree distribution. We discuss this further in Section 6.
@@
Similarly, though the spanning tree distribution has a close connection to random walks on the underlying graph, a significant amount of the randomness is “lost” when constraints such as balance are added. For example, consider the probability that two vertices are separated in the spanning tree distribution $\mathcal{D}_{\mathtt{sp}}$. For $k=2$, this probability can be shown to be directly proportional to the effective resistance between the two vertices barrett2019spanning2forestsresistancedistance ; wu15-sepprob ; wu23-sepprob . On the other hand, no general expression is known for this probability even under mild modifications of the distribution towards balance.
@@
Definition 9 and Theorem 6 form the heart of our technical contributions. The former and the associated Theorem 7 shows that our modified process indeed allows us to analyze partitions purely from the perspective of random cycles. Our main contribution here is a constructive mapping between random walks: given a walk whose erasure separates two vertices, we devise an intricate algorithm that builds a corresponding walk that does not, carefully re-splicing all the original loops to ensure the mapping is one-to-one, and the number of different directed edges is small. This formalizes the intuition that a local change to a partition’s boundary should correspond to a local change in the underlying random walk. This result can be viewed as a complement to the main result from procaccia2022compact - they show that two partitions whose boundaries greatly differ (in terms of length) must have greatly different probabilities of being sampled whereas ours shows that two partitions with very similar boundaries have similar probabilities of being sampled. The latter result, Theorem 6, proves that such a local modification is always possible. To do this, we develop a suite of graph-theoretic tools in Section 5.2 specifically for analyzing 2-partitions on grid graphs. Our key insight is the following: After switching vertices from one part to the other, either the original partitions were both “thick” and this did not disconnect any part, or they were “thin” and there must exist some structures that we can identify and perform local surgery on (Lemmas 25 and 26). Through an extensive case analysis facilitated by these tools, we prove that any separating partition can be modified in a small neighborhood to become non-separating, with only a constant change to the imbalance.
@@ 3.1 Modified Random Walk Algorithm
where the sum is over all connected partitions $X,Y$. Similarly, for any pair of vertices $u,v\in V(G)$, let $\texttt{sp}(G,2)_{u\not\sim v}$ denote the same sum except the sum is over all connected partitions $X,Y$ such that $u\in X$, $v\in Y$. We note that these normalization factors do not need to be computed, since we do not intend to actually run our algorithm, only analyze it. [^5] Nevertheless, the required factor can be computed since $\texttt{sp}(G,2)_{u\not\sim v}$ is simply the number of spanning trees of $G$ containing the edge $(u,v)$. For our purpose, we can replace $\texttt{sp}(G,2)$ in the denominator with the total number of spanning trees instead. The ratio of $\texttt{sp}(G,2)_{u\not\sim v}$ and the number of spanning trees is the effective resistance between $u,v$ barrett2019spanning2forestsresistancedistance , which can be computed in polynomial time. We are now ready to state Algorithm 2.
@@ 3.3 Unseparating Mappings imply Theorem 1
We assume Theorems 6 and 7 are true. Claim 4 implies that that we can sample from $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ as follows: First, run Algorithm 2. Accept the resulting partition $P$ with probability $e^{-\lambda\cdot\texttt{imb}(P)}$ and restart otherwise. Fix some pair of adjacent vertices $u,v$ and let $f$ be the promised unseparating mapping for them. Let $H^{*}$ be the subgrid in Property 1.
@@
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}\wedge\mathbf{E}_{2\beta}]\geq\frac{1}{3\beta e^{\delta\lambda}4^{3\beta^{2}}}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}\wedge\mathbf{E}_{2\beta}].$ (4)
@@
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}\wedge\mathbf{E}_{2\beta}]$ $\geq\frac{1}{3\beta e^{\delta\lambda}4^{3\beta^{2}}}(\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}]-\frac{1}{2})$
$\implies\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are not separated}]$ $\geq\frac{1}{3\beta e^{\delta\lambda}4^{3\beta^{2}}}(\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[u,v\text{ are separated}]-\frac{1}{2})$
@@
3.4 Extension to Groups of Events
In this section, we extend Theorem 1 by showing that the probability that multiple edges are separated in a partition sampled from $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ is also small. We follow an argument similar to the proof of Theorem 1, except we apply the unseparating mapping multiple times. More specifically, instead of comparing the probability that an edge $e$ is separated with the probability that it is not separated, we will compare the probability that $q$ edges are simultaneously separated and the probability that only a subset of them are separated.
Theorem 9.
For constant $\lambda$, and $q=o(\sqrt{mn})$, and any set of $q$ edges of $\boxplus_{m\times n}$, each distance at least $\gamma+1$ apart (where $\gamma$ is the constant from Theorem 21), the probability that a 2-partition sampled from $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ separates all the $q$ edges is $e^{-qc}$, for some constant $c$.
Proof.
As before, we will assume the correctness of Theorems 6 and 7. Fix some set $Q$ of $q$ edges that are each at distance at least $\gamma+1$ from each other. Let $\mathcal{E}$ be an arbitrary subset of $Q$ of size $\ell$ (to be determined later). Order the edges of $\mathcal{E}$ as $e_{1},\ldots,e_{\ell}$. Let $f_{1},f_{2},\ldots,f_{\ell}$ be the unseparating mappings for them.
Consider some cycle $C$ output by Algorithm 2 that separates every edge in $\mathcal{E}$. We assume that $C$ has length at least $2\ell\beta$. Define a set of functions $F_{i}$ such that $F_{0}(C)=C$ and $F_{i}(C)=f_{i}(F_{i-1}(C))$ for all $1\leq i\leq\ell$. Since each unseparating mapping can decrease the length of the cycle by at most $\beta$ (because of Property 1), this is well-defined. Since the subset $\mathcal{E}$ and an arbitrary ordering of the initial $q$ edges uniquely defines $F_{\ell}(C)$, we will also refer to $F_{\ell}(C)$ as $F(\mathcal{E},C)$. Define $H_{i}$ to be the subgrid (from Property 1) where the cycles $F_{i}(C)$ and $F_{i-1}(C)$ differ. Note that since the edges of $\mathcal{E}$ are at distance at least $\gamma+1$ from each other, the unseparating mappings have the additional property that the graphs $H_{i}$ are disjoint.
[…]
Like before, we now apply Theorem 7 to $F_{i}(C)$ and $F_{i+1}(C)$, sequentially for $i=0,1,\ldots,q$, so that for any walk $\Pi$ whose loop erasure was $C$, we get a unique walk $\Pi^{\prime}$ whose loop erasure is $F(\mathcal{E},C)$, such that $\Pi^{\prime}$ has at most $3\ell\beta^{2}$ additional edges that were not in $\Pi$, and none of these edges were incident to $v_{0}^{*}$. Repeating the same arguments as before, across all possible random walks $\Pi$, and noting that the population imbalance can increase by at most $\ell\delta$, we get
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[F(\mathcal{E},C)\text{ is sampled}]\geq\frac{1}{3e^{\ell\delta\lambda}4^{3\ell\beta^{2}}}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[C\text{ is sampled}]$ (5)
We now want to add up the above equation over different cycles $C$ and subsets $\mathcal{E}$ of $Q$. To this end, consider when $F(\mathcal{E},C)=F(\mathcal{E}^{\prime},C^{\prime})$ for some other subset $\mathcal{E}^{\prime}\subseteq Q$ of size $\ell$ and some other cycle $C^{\prime}$ that separates all the edges in $Q$. Firstly, since every edge from $Q$ except those in $\mathcal{E}$ are still separated in $F(\mathcal{E},C)$, and both $\mathcal{E}$ and $\mathcal{E}^{\prime}$ have size $\ell$, it follows that $\mathcal{E}=\mathcal{E}^{\prime}$. Thus, the only cycles $C^{\prime}$ for which $F(\mathcal{E},C)=F(\mathcal{E},C^{\prime})$ is when the unseparating mapping maps two cycles that separate one edge in $\mathcal{E}$ to the same cycle that does not separate it. From Property (4) of unseparating mappings, there are at most $\beta$ such cycles per edge in $\mathcal{E}$, meaning that there are at most $\beta^{\ell}$ cycles $C^{\prime}$ such that $F(\mathcal{E},C)=F(\mathcal{E},C^{\prime})$.
Recall that we assumed that $C$ has length at least $2\ell\beta$. Let $\mathbf{E}_{2\ell\beta}$ denote the event that $\lambda$-$\mathcal{D}_{\mathtt{sp}}$ samples a cycle of length at least $2\ell\beta$. Adding up Eq. 5 over all cycles $C$ of length at least $2\ell\beta$ that separate every edge in $Q$, and over all subsets $\mathcal{E}$ of $Q$ of size $\ell$, we get
$\sum_{\mathcal{E}}\sum_{C}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[F(\mathcal{E},C)\text{ is sampled}]$ $\geq\binom{q}{\ell}\sum_{C}\frac{1}{3e^{\ell\delta\lambda}4^{3\ell\beta^{2}}}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[C\text{ is sampled}]$
The RHS is precisely the probability that every edge in $Q$ is separated intersected with the event $\mathbf{E}_{2\ell\beta}$. The LHS is a lower bound on the complementary event (i.e. not every edge in $Q$ is separated), except we may be over counting each cycle $\beta^{\ell}$ times. We can bound this above by $\beta^{\ell}$. Thus,
$\beta^{\ell}$ $\geq\binom{q}{\ell}\frac{1}{3e^{\ell\delta\lambda}4^{3\ell\beta^{2}}}\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[\text{Every edge in }Q\text{ is separated}\wedge\mathbf{E}_{2\ell\beta}]$ (6)
Next, claim 8 implies that a cycle with length smaller than $2\ell\beta$ encloses at most $12\ell^{2}\beta^{2}$ vertices, implying $\texttt{imb}(C)\geq\frac{mn}{2}-12\ell^{2}\beta^{2}$. Since $\beta,\lambda$ are constants, for large enough $m,n$ and $q=o(\sqrt{mn})$, we can argue similarly as in Appendix A that $\Pr[\neg\mathbf{E}_{2\ell\beta}]\leq\frac{1}{2}e^{-qc}$ for some constant $c$. Then, since we have
$\Pr[\text{Every edge in }Q\text{ is separated}\wedge\mathbf{E}_{2\ell\beta}]$ $\geq\Pr[\text{Every edge in }Q\text{ is separated}]-\Pr[\neg\mathbf{E}_{2\ell\beta}],$
it must hold that either $\Pr[\text{Every edge in }Q\text{ is separated}]\leq e^{-qc}$ is true, in which case the proof is done, or it is the case that
$\Pr[\text{Every edge in }Q\text{ is separated}\wedge\mathbf{E}_{2\ell\beta}]$ $\geq\frac{1}{2}\Pr[\text{Every edge in }Q\text{ is separated}].$
Plugging this back into Eq. 6, we get
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[\text{Every edge in }Q\text{ is separated}]$ $\leq\frac{6}{\binom{q}{\ell}}\left(\beta e^{\delta\lambda}4^{3\beta^{2}}\right)^{\ell}.$
Since $\beta,\lambda$ are constants, $c=\beta e^{\delta\lambda}4^{3\beta^{2}}$ is also a constant. We lower bound $\binom{q}{\ell}$ with $\left(\frac{q}{\ell}\right)^{\ell}$. The expression is then minimized at $\ell=\frac{q}{ec}$, giving us
$\Pr_{\lambda-\mathcal{D}_{\mathtt{sp}}}[\text{Every edge in }Q\text{ is separated}]$ $\leq 6e^{-\frac{q}{ec}}$
as required. ∎
@@ 4.2 Proof of Theorem 7
Now, recall that Algorithm 3 added these walks to $W^{\prime}$ in such a way that $N_{1}$ appears first in $\widehat{W}$, $N_{2}$ appears just after that, and so on, so that $N_{p}$ appears last. Let $L_{1},L_{2},\ldots,L_{m^{\prime}}$ be the loops added because of $N_{1}$. That is, $N_{1}=L_{1}L_{2}\ldots L_{m^{\prime}}$. We want to show that in the first iteration of step 12 in Algorithm 4, we correctly identified $N_{1}$. For any $1\leq p\leq m^{\prime}$, let $w_{p}$ be the smallest element in $L_{p}$ according to $\preceq_{D}$, so that $\{w_{1},w_{2},\ldots,w_{m^{\prime}}\}$ is the loop sequence of these loops. Since each $L_{p}$ (or possibly a cyclic permutation of it) was a part of $M_{1}$, which was a maximal loop at $v_{i}$ in $W$, we can apply Observation 12 to get that either $w_{p}=\varphi$ or $v_{1}\preceq_{D}w_{p}$. Further, since $N_{1}$ must be the cyclic permutation of $M_{1}$ from the first appearance of $v_{1}$, $v_{1}$ must appear on $L_{m^{\prime}}$. If $v_{1}$ did not appear on $L_{m^{\prime}}$ but only appeared on, say, $L_{m^{\prime}-1}$ instead, this would mean that $N_{1}$ is the cyclic permutation of $M_{1}$ from the second appearance of $u_{i}$. Thus, $w_{m^{\prime}}=v_{1}$. We now argue that $m^{\prime}$ is the first element of the breakdown of the loop-sequence of $\mathcal{L}$. This will complete our argument that $N_{1}$ was correctly identified in step 12. To see this, observe that by construction of $\mathcal{M}_{u_{i}}$, $v_{1}\prec_{D}v_{2}\prec_{D}\ldots\prec_{D}v_{m}$. Further, the walks $M_{2},M_{3},\ldots,$ were comprised of maximal loops from $v_{2},v_{3},\ldots$ respectively. From Observation 12, for any $p>m^{\prime}$, it must be that $w_{m^{\prime}}=v_{1}\prec_{D}w_{p}$, or $w_{p}=\varphi$. Combined with our earlier observation that $0<p<m^{\prime}\implies v_{1}\preceq_{D}w_{p}$, or $w_{p}=\varphi$, this means that $v_{1}$ is in the breakdown of the loop-sequence of $\mathcal{L}$. It also must be the first element of the breakdown, since if there exists some $w_{p}$ with $p<m^{\prime}$ and $w_{p}\prec_{D}v_{1}$, this creates a contradiction with our observations. Repeating the arguments for $N_{2},N_{3},\ldots,$ in the same way completes the proof of Subclaim 0.1, when $\mathcal{M}_{u_{i^{\prime}}}=\phi$ for all $i^{\prime}<i$.
@@
This concludes the proof of Lemma 17. Along with Claims 14 and 16, this also completes the proof of Theorem 7.